# A Clinical Tuning Framework for Continuous Kinematic and Impedance Control of a Powered Knee-Ankle Prosthesis

### v1

@@ III Results
Figure 5: Knee (top) and ankle (bottom) torque from the prosthesis at each incline in the baseline (solid) and individualized (dashed) controllers. The shaded regions indicate standard deviation across all strides.
@@
Figure 7: Knee Moments for Sit-to-Stand (top) and Stand-to-Sit (bottom) in the tuned and baseline controllers. The motions are aligned around peak moment and are measured in frames (250Hz). The mean $\pm$ SD is shown in dashed lines.
@@
Next, we look at the ankle push-off tuning in the CTI. Figure 5 shows that while a +50% change was applied in the tuning parameter, the achieved peak torques increased closer to 25% across tasks (20%, 24%, and 29% at decline, level, and inclined tasks). Each incline shows significant increases, but not of the magnitude expected. Looking at the impedance parameters in Figure 4, we can see that the optimization accommodated the request for additional ankle-pushoff by increasing damping in late stance and increasing stiffness throughout the gait cycle. The ankle kinematics in Figure 6 exhibit a slight shift in phase on the prosthetic side around the plantarflexion peak, though no major changes in magnitude can be observed from the added push-off torque. On the sound side, there is a significant shift in both magnitude and phase of motion during late stance/early swing (25-75%). Similar changes in phase are observed in declined and level walking, though not as extreme as inclined walking. The change in shape around 40% gait is more similar to the prosthetic side, indicating increased inter-limb symmetry.
While no changes in peak knee flexion were prescribed in the swing kinematic model, we observe differences in peak knee flexion in practice during inclined and declined walking, 1.4° and 2.6° respectively (Figure 6). On the prosthetic side, a change occurs in the shape of the first peak of knee flexion that corresponds with the loading of a stiffer knee. On the sound side, we observe increases in knee flexion at all inclines, as well as a shift in phase of both peak knee flexion and extension that aligns better with the prosthetic side, indicating a more symmetrical gait. Further, while the CTI did not allow changes to the swing ankle kinematics, we did see a slight change on the prosthetic side that we attribute to the increased push-off moment.
@@
IV Discussion and Conclusion
TF01’s extensive experience with the un-tuned controllers led to mixed preferences. The subject selected significant changes to the controller and could easily perceive the differences during tuning sessions. However, when blinded to conditions, TF01 preferred the baseline walking controller. For sit-to-stand transitions, TF01 favored increased assistance when rising from a seated position but chose lower sitting torque to enable faster motion. The parameter distinction between sitting and standing was crucial for successful controller tuning. We observed that more extreme modifications led to the exclusion of trajectories from the training set, indicating increased difficulty in solving the problem. Nonetheless, the accepted models maintained errors below 5% at the ankle and 15% at the knee, effectively reducing commanded peak values for tasks.
@@ IV Discussion and Conclusion
There are a few points that we found interesting in this work. Firstly, the baseline controller is designed to emulate normative able-bodied motion and the subject chose strong parameter changes during tuning, but during walking trials TF01 indicated a preference for the normative, biomimetic motion. Secondly, knee swing flexion had complex effects on gait, balancing toe clearance with shank progression speed for comfortable socket-limb dynamics. Lastly, tuning specific parameters led to cascading effects on gait, especially as tuning diverged from normative values, suggesting avenues for future investigation.
@@
This study has several limitations with potential implications. Firstly, the typical approach of refining tuning over repeated clinical sessions, as done with commercial prostheses, was not feasible due to constraints of the research hardware. Secondly, a larger subject sample size is necessary to comprehend the range of parameter changes in the population and associated discrepancies between desired and realized torques. Finally, observational gait analysis is inherently subjective, and tuning outcomes rely on both clinician and subject feedback.
The definition of the IKC can be found in the methods section, and the IKC trajectories of a representative subject across modes and tasks are shown in Figure 1.
-A Individualization of Walking Kinetics
@@
-B Kinetic Individuality Validation
The results of the statistical validation of kinetic individuality based on the offline dataset data are as follows. The baseline method was more predictive than the LOO for 76% of the all subjects and joints, showing significant improvements upon individualization for each joint across all tasks. There was a 44% (0.085 Nm/kg) improvement in fit at the ankle, 18% (0.024 Nm/kg) at the knee, and 31% (0.05 Nm/kg) at the hip (Figure 8). Each of the joints shows a significantly improved fit upon individualization with initial one-tailed paired t-test values of $p=0.002$, $p=0.011$, and $p=0.003$ for the ankle, knee, and hip, respectively. Subject and joint specific details are given in Table I.
Figure 8: Kinetic RMSE improvement (positive) over LOO average for walking tasks at 1.0m/s for the ankle (green), knee (blue), and hip (orange). The boxes show the 25-75% interquartile range, with a line at the median; the whiskers show the non-outlier maximums. Each bar represents the data across all subjects.
@@
Table I: This table shows RMSE of each predicted trajectory when compared to the experimental values, similar to the kinematic calculation seen in (9). Significance between walking and LOO values is represented with ‘*’ for $p<0.05$. These values are the average error across the gait cycle in Nm/kg.
@@ Appendix A Discussion
A-A Kinetic Individuality Conclusions
Our method for predicting kinetic individuality across tasks was more accurate than the LOO average alone (Figure 8). We observed more pronounced trends within subjects regarding kinetic individuality compared to those reported for kinematics in [24], with significant improvements for all joints, when all subjects, tasks, and strides are considered (see Table I).
@@ A-A Kinetic Individuality Conclusions
As discussed earlier, this method of propagating individuality had mixed results across subjects, indicating that baseline individuality is more predictive across tasks for some subjects than others (Table I). For example, AB03 shows good improvement at all joints, while AB07 sees the opposite trends. These trends across subjects demonstrate particularly accurate prediction when the LOO average kinetics are different from the population average, as observed with AB04. This is especially encouraging as we are applying this method to users who have impaired walking.

### v2

@@ 3 Results
Figure 6: Prosthetic knee (top) and ankle (bottom) torques at each incline in the baseline (solid) and individualized (dashed) controllers. The shaded regions indicate standard deviation across all strides.
@@
Figure 8: Prosthetic knee torques for Sit-to-Stand (top) and Stand-to-Sit (bottom) with the tuned and baseline controllers. Solid lines represent the mean and dashed lines represent $\pm$ SD.
@@
Next, we look at the ankle push-off tuning in the CTI. Figure 6 shows that while a +50% change was applied in the tuning parameter, the achieved peak torques increased closer to 25% across tasks (20%, 24%, and 29% at decline, level, and inclined tasks). Each incline shows meaningful increases, but not of the magnitude expected. Looking at the impedance parameters in Figure 5, we can see that the optimization accommodated the request for additional ankle pushoff by increasing stiffness throughout the gait cycle and increasing damping in late stance. The prosthetic ankle kinematics in Figure 7 exhibit a slight shift in phase around the plantarflexion peak, though no major changes in magnitude can be observed from the added push-off torque. On the sound side, there is a significant shift in both magnitude and phase of motion during late stance/early swing (25-75%). Similar changes in phase are observed in declined and level walking, though not as extreme as inclined walking. The change in shape around 40% gait is more similar to the prosthetic side, indicating increased inter-limb symmetry.
Figure 7 also reveals changes in the shape of the first peak of prosthetic knee flexion that corresponds with loading a stiffer knee. On the sound side, we observe increases in knee flexion at all inclines, as well as a shift in phase of both peak knee flexion and extension that aligns better with the prosthetic side, indicating a more symmetrical gait.
@@
4 Discussion and Conclusion
TF01’s extensive experience with the un-tuned controllers led to mixed preferences. The subject selected significant changes to the controller and could easily perceive the differences during tuning sessions. However, when blinded to conditions, TF01 preferred the baseline walking controller. We believe that this is largely based on the subject’s prior experience with the baseline controller and inadequate acclimation time with the tuned controller. For sit-to-stand transitions, TF01 favored increased assistance when rising from a seated position but chose lower sitting torque to enable faster motion. The parameter distinction between sitting and standing was crucial for successful controller tuning. We observed that more extreme modifications led to the exclusion of trajectories from the training set, indicating increased difficulty in solving the problem. Nonetheless, the accepted models maintained errors below 5% at the ankle and 15% at the knee, effectively reducing commanded peak values for tasks.
@@ 4 Discussion and Conclusion
There are a few points that we found interesting in this work. Firstly, the baseline controller is designed to emulate normative able-bodied motion and the subject chose strong parameter changes during tuning, but during walking trials TF01 indicated a preference for the normative, biomimetic motion. Secondly, knee swing flexion had complex effects on gait, balancing toe clearance with shank progression speed for comfortable socket-limb dynamics. Lastly, tuning specific parameters led to cascading effects on gait, especially as tuning diverged from normative values. These findings highlight potential directions for future research and the development of automated tools to enhance the efficiency and effectiveness of the tuning process.
@@
Acknowledgment
This work was supported by the National Institute of Child Health & Human Development of the NIH under Award Number R01HD09477. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
This appendix details the validation of the individual kinetic contribution (IKC, Eq. 2). The IKC of a representative subject across modes and tasks are shown in Figure 2.
Individualization of Walking Kinetics
@@ Acknowledgment
Kinetic Individuality Validation
The results of the statistical validation of kinetic individuality based on the offline dataset data are as follows. The baseline method was more predictive than the LOO for 76% of the all subjects and joints, showing significant improvements upon individualization for each joint across all tasks. There was a 44% (0.085 Nm/kg) improvement in fit at the ankle, 18% (0.024 Nm/kg) at the knee, and 31% (0.05 Nm/kg) at the hip (Figure 9). Each of the joints shows a significantly improved fit upon individualization with initial one-tailed paired t-test values of $p=0.002$, $p=0.011$, and $p=0.003$ for the ankle, knee, and hip, respectively. Subject and joint specific details are given in Table 1.
Figure 9: Kinetic RMSE improvement (positive) over LOO average for walking tasks at 1.0m/s for the ankle (green), knee (blue), and hip (orange). The boxes show the 25-75% interquartile range, with a line at the median; the whiskers show the non-outlier maximums and circles are outliers. Each box represents the data across all subjects.
@@
Table 1: RMSE of each predicted trajectory when compared to the experimental values, similar to the kinematic calculation seen in (9). Significance between walking and LOO values is represented with ‘*’ for $p<0.05$. These values are the average error across the gait cycle in Nm/kg.
@@ Appendix A Discussion
Kinetic Individuality Conclusions
Our method for predicting kinetic individuality across tasks was more accurate than the LOO average alone (Figure 9). We observed more pronounced trends within subjects regarding kinetic individuality compared to those reported for kinematics in [25], with significant improvements for all joints, when all subjects, tasks, and strides are considered (see Table 1).
@@
As discussed earlier, this method of propagating individuality had mixed results across subjects, indicating that baseline individuality is more predictive across tasks for some subjects than others (Table 1). For example, AB03 shows good improvement at all joints, while AB07 sees the opposite trends. These trends across subjects demonstrate particularly accurate prediction when the LOO average kinetics are different from the population average, as observed with AB04. This is especially encouraging as we are applying this method to users who have impaired walking.